### v2

# Shavette: Safe Undervolting of Neural Network Accelerators via Algorithm-level Error Detection

Abstract.
With transistor technology reaching its limits, exploring alternative approaches to enhance the energy efficiency of Deep Neural Network (DNN) accelerators has become essential. Operating at lower voltages substantially enhances energy efficiency. However, this approach is limited by increased sensitivity to Process, Voltage, and Temperature (PVT) variations, which complicates identifying the optimal energy-efficient voltage-frequency operating point under throughput constraints. Traditional approaches to adaptive voltage scaling significantly increase design complexity and are unsuitable for already fabricated devices.
This paper introduces a novel application of algorithmic fault detection methods to enable reliable and aggressive voltage down-scaling in DNN accelerators. While error detection approaches have been widely studied in the context of system reliability, this work is the first to leverage these techniques specifically for reducing operational voltage in DNN accelerators as a practical, low-cost energy-saving solution. Two popular DNN models, i.e., LeNet and VGG-16, were developed from scratch in C++/OpenCL to integrate the algorithmic error protection mechanisms. Experiments on a Graphical Processing Unit (GPU) demonstrated $18\%$ to $25\%$ energy savings with no accuracy loss and trivial, $<4\%$, throughput compromise.
Low power, NPU, Voltage Scaling, DNN Accelerators, ABFT, Energy Efficiency
1. Introduction
Deep Neural Networks (DNNs) are becoming increasingly popular for consumer applications, including 5G connected edge devices such as autonomous vehicles (Tarkoma et al., 2023). The reliance on DNNs in these devices presents a growing challenge. Their computational demands are substantially higher than conventional embedded applications, hence, DNNs strongly impact energy efficiency, straining battery life and impose severe thermal constraints. Already being at the edge of transistor technology, i.e., due to end of Moore’s law and Dennard scaling, utility of alternative energy optimization approaches becomes more prominent.
One such approach involves operating at reduced voltages (Krishna, 2022). Lowering voltage quadratically reduces power consumption, e.g., a mere 20% voltage reduction can cut the energy consumption to half. In the vicinity of the threshold voltages of the transistors, $V_{th}$, energy efficiency gains of 10x to 100x are provisioned (Ernst et al., 2003). Unfortunately, as the operating voltage is reduced, sensitivity to Process, Voltage and Temperature (PVT) variations exacerbates, rendering commonly used Static Timing Analysis (STA) methods unreliable.
Voltage reduction necessarily requires lowering frequency to meet timings (Ernst et al., 2003). However, manufacturers often include a safety margin, typically around 10% to 20%, in their recommended voltage, implying the device might function with the same performance at even lower voltages. This margin accounts for potential variations in temperature, voltage, and process conditions. Studies have found different margin values: 10% in commercial CPUs (Krishna, 2022), 20% in Field-Programmable Gate Arrays (FPGAs) (Safarpour et al., 2021b), and around 10% in GPUs (Thomas et al., 2016). If those voltage margins are eliminated safely, substantial power savings without throughput lost, i.e., frequency down-scaling, is achievable (Thomas et al., 2016). It is difficult to remove margins in a manner that ensures no timing errors in logic circuits or bit-flip in memory systems are induced. The margin size varies due to PVT variations from die to die and as a function of time for the same chip, i.e., due to aging. STA does not provide reliable operating points either, hence, modeling the processor for optimal operating point, that is the lowest working voltage and highest clock frequency, becomes very challenging. GPUs in particular experience larger voltage variation due to IR drop, which needs to be considered (Thomas et al., 2016).
Traditionally, voltage margin trimming techniques, e.g., Razor (Ernst et al., 2003), rely on adding extra hardware into the digital circuit. Those were not conventionally supported by Computer Aided Design (CAD) tools and add substantial design complexity and incur significant increase of testing costs (Hiienkari et al., 2020). In contrast, this contribution introduces a straightforward, low-overhead solution [^1] Source codes associated with this paper are available online at https://github.com/NortHund/OCDNN. to enable reliable elimination of voltage margins in DNN accelerators.
To the best knowledge of authors, this is the first work on leveraging fault detection techniques for enabling a low-cost fully software undervolting solution. The work demonstrates the solution using real-world DNN model and processing platform. The results show substantial savings.
2. Background and Motivation
Two of the most investigated hardware based techniques for enabling reduced voltage operation are briefly discussed here. A simple option is to use adaptive delay monitoring circuits, e.g., via ring oscillators or delay chains. This solution offers wider applicability but suffers from unreliability and does not ensure optimal voltage-frequency estimates (Kim et al., 2016). The second approach is based on inserting Timing Error Detection (TED) based systems (Ernst et al., 2003), like Error Detection Sequences (EDS), as is depicted in Fig. 1, into the critical paths of the circuit to monitor for any timing errors in-situ. As shown in Fig. 1, a set of extra secondary flip-flops clocked with a slight delay compared to the primary set of flip-flops are inserted the end of critical combinational logic paths. By detecting a late arrival signal, possibly due to reduced voltage operation, the timing errors become detectable. The detections from TED system are fed back into the voltage regulator to adjust the voltage such that minimum voltage is selected while no errors are observed. Although TED based solutions, e.g., the well-known Razor (Ernst et al., 2003), enable significant voltage reduction, they come with high development complexities and overheads as those techniques cause significant hiccups in semiconductor design and manufacturing flow, e.g., by requiring extensive cell library characterization or secondary clock tree insertion. Finally, the hardware based techniques are only implementable in pre-silicon stage (do Nascimento et al., 2023).
Figure 1. Simplified concept of Timing-Error-Detection circuit (Safarpour et al., 2021a).
3. Proposed Solution: Algorithmic Error Detection
As mentioned earlier, for enabling error-free computations at reduced voltages, a run-time error detection mechanism is needed. As an alternative to hardware based methods for enabling low voltage operation, the proposed solution relies on Algorithm Based Fault Tolerance (ABFT) which was originally introduced by Huang and Abraham (Huang and Abraham, 1984) in 80s for detection and correction of errors in large matrix computations.
3.1. Detecting error in linear layers
[…]
3.2. Detecting errors in non-linear operations
Since the ABFT schemes only work for linear operations while DNNs consist of non-linear layers, e.g., pooling and activation functions (LeCun et al., 1998), Double Redundant Module (DMR) approach was employed for those layers. Although DMR incurs large overheads, i.e. 100%, but fortunately, computations of non-linear layers contribute to just of a few percent of total computations (LeCun et al., 1998). Hence, error detection through DMR adds tolerable overheads, when is restricted to only non-linear layers. Note, the redundant module must be implemented uncorrelated to the original one, e.g., with different instruction set. This is to avoid correlated errors. In experiments it was “observed” that errors always appear in linear operations before non-linear operations possibly due to exercising longer delay paths. Nonetheless DMR was still added to protect all layers against errors (Nia and Mohammadi, 2007).
Notice, although ABFT and DMR have been existing for long, to the best knowledge of authors, this is the first time they are leveraged for reducing voltage/power of DNN processing.
Figure 2. Checksums enable detections of computing errors (adapted from (Safarpour et al., 2021a)).
4. Experiments and Results
In our implementation the inference of ABFT augmented DNN model is carried out within the undervolted accelerator, i.e., GPU, while its voltage is controlled by the CPU. The inference output, i.e., DNN’s prediction, along with checksums and DMR results are sent back to the CPU. The CPU checks for computational errors by verifying the correctness of linear and non-linear operations within the model as described in previous subsections. If no errors are observed the inference results from the neural network model is accepted, otherwise, e.g., the voltage/frequency is adjusted and inference operation is repeated.
In this manner, the voltage of the GPU as the DNN accelerator, is dynamically adjusted to minimum error free operating voltage, without incurring errors in computations, neither degrading network accuracy nor causing system crash by timely detection of the Point of First Failure (PoFF). No additional on-chip or off-chip modifications are involved and hence the solution is truly applicable to the commercial off-the-shelf components (Rinkinen, 2024).
To detect errors in linear parts, i.e., the convolutional and fully-connected layers, the convolution and matrix operations kernels were modified. Details of implementation are available in our online Git repository mentioned earlier. With non-linear layers the DMR modules used two different codes to avoid the same faults appearing in both, causing incorrect voting results.
ABFT checksum for each layer was calculated on-the-fly within the GPU. The checksums of weights were pre-computed for each model to minimize overheads. For inference, pre-computing the input checksums for weights is possible, however, training obviously requires updating the weights and hence re-computing the wight checksums.
The scheme in our experimentation was to keep the clock frequency and reduce the operating voltage until errors start appearing. Once errors detected the voltage can be retracted to slightly higher voltage. However, we continued reducing voltage to detect both the PoFF and point of crash to assess the error detection capability. PoFF voltage turned out to be much higher than crash voltage, hence, the solution alarms voltage governor, way before approaching the crash point. In all experiments, the on-chip temperature sensor report was fluctuating around $55^{\circ}C$. Notice at each voltage step same test batch was used for estimating the accuracy.
Figure 3. Power consumption with ABFT enabled and ABFT disabled for both VGG-16 at 1780 MHz and 1680 MHz clock frequencies with respect to voltage. The PoFF and crash points are marked for each.
4.1. Power Saving
A particularly interesting issue is the dynamic power savings of the GPU. As the voltage scaling quadratically reduces dynamic power, we kept highest utilization possible and scaled down the voltage till an error is observed by ABFT, i.e., PoFF is detected. Once PoFF is determined, one could stop reducing voltage; however, for characterization purposes we further reduced voltage down to the crash point. At the same time, the actual errors and impact of them on accuracy loss of the DNN model was measured. The bar, i.e., error threshold, for reporting using ABFT was set high which results in the higher error rate reported by ABFT compared to actual error rate. One could set a more relaxed error threshold, adding few more milliwatts of power savings. Figure 3 depicts power savings trends while Fig. 4 plots the error rate and accuracy loss for different voltages. Notice, the PoFF and crash points changes for different clock frequencies.
We initially implemented ABFT on LeNet to assess its viability. However, due to the small model size of LeNet and the inherent overhead scaling of ABFT, which is $(1/N)$, where $N$, represents the problem size, we found that ABFT is not well-suited for very small deep neural networks (DNNs). While the preliminary results demonstrated notable power savings, the observed 7% overhead could be misleading and may lead to incorrect generalizations about ABFT’s efficiency. In contrast, VGG-16, with its significantly larger size, exhibited much lower relative overheads, making it a more appropriate candidate for evaluating ABFT’s impact on larger DNNs. Hence, LeNet results were not plotted. Nonetheles, those are reported here. The power drops from around 30W down to 25W at 812mV (PoFF for LeNet) and 22W at around 805mV and $f_{clk}=1780MHz$.
Measurements results for power are averaged out of 120 inference experiments per operating points taken within 0.5s intervals. Table  1 shows the power dissipation differences of the GPU at nominal voltage $V_{default}$ and three PoFF voltages $V_{min}$ and respective clock rates along with calculated energy overheads.
Table 1. Power savings and energy overheads of ABFT-enabled VGG-16 at different clock frequencies.
Parameter 1820 MHz 1780 MHz 1680 MHz
P($W$) @Nominal $141$ $142$ $137$
$V_{min}$ (mV) $850$ $835$ $800$
P($W$) @$V_{min}$ $116$ $110$ $107$
Energy/Inference ($J$) $20.7$ $19.9$ $19.0$
[…]
Figure 4. The voltage is reduced from detail down to the crash point. The ABFT detect errors in computations. Notice no accuracy loss is observed despite ABFT detections which is due to inherent fault tolerance of DNNs.
4.2. Error Coverage
Figure 4, depicts the errors appearing in computations, i.e., actual error rate, error rate from ABFT detections and overall prediction errors (accuracy loss) of the model. Since errors propagate from one layer to another, comparing ABFT error detection rate against actual errors per each operations is misleading. Experiments on single operation, shows very high (close to 100%) computational detection rate for ABFT, similiar to results on in (Safarpour et al., 2021b). Errors were reported if difference between the GPU accelerated outputs and the reference was over 1e-13. The error threshold can be made slightly tighter to ensure more reliability, however, that would result in false positives being detected constantly, even at stock voltage and frequency settings.The DNN accuracy was compared against baseline (not undervolted). Notably, due to inherent fault-tolerance of DNNs a few hundred errors at vicinity of PoFF does not impact the accuracy, as it is observed in Fig. 4. Nonetheless, since ABFT detects the computational errors early enough, we do not need to rely on fault-tolerance feature of DNNs for ensuring reliability.
As expected the errors appear in linear layers significantly before being detected in the non-linear ones. We speculate the possible explanation is that the circuit delay paths for non-linear operations are much shorter than those of floating-point multipliers used for arithmetic computations.
4.3. Overheads
The power consumption at all voltages with ABFT enabled is slightly lower (around 1%) than when it is disabled, meaning ABFT overheads manifest in inference time rather than power. That is likely due to ABFT related calculations imposing more idle periods on processing units. As Table  1 shows, this translates into slightly longer inference time. Overall energy overhead estimated by comparing the energy per inferences between ABFT enabled and disabled runs. Table 2 provides a summary of overheads in terms of inference time.
Naive integration of ABFT into the DNN model results in large overheads, however, with small tweaks and code optimization the overhead share becomes insignificant, e.g., $<3.9\%$ for VGG-16. Considering VGG-16 with default clock frequency and voltage (at 1780 MHz and 960 mV), the energy overhead of ABFT is as low as 1%. With same frequency but voltage set to 880 mV, the energy overhead is almost the same, i.e., 1.3%. Setting the frequency to the lowest, i.e., 1680 MHz, and at the lowest stable around 805 mV, the energy overhead of ABFT enabled is slightly higher, i.e., 1.9%. This is a small penalty for saving almost a quarter of the energy of each inference.
Table 2. Inference times VGG-16 at different clock frequencies.
Configuration VGG-16 VGG-16 VGG-16
@ 1820 MHz @ 1780 MHz @ 1680 MHz
ABFT disabled ($ms)$ 171.90 175.19 183.42
ABFT enabled ($ms$) 178.08 181.36 189.86
Time overhead (%) $\approx 3.5\%$
5. Related works
In (Gundi et al., 2020), a Tensor Processing Unit (TPU) architecture, named ”EFFORT”, is proposed. The TPU is equipped with timing error detection circuits within critical paths of Multiply and Accumulate (MAC) units to enable reduced voltage operations. EFFORT achieves high performance and energy efficiency by employing TED systems and in-situ clock gating techniques to address. To minimize hardware overheads, not all critical paths can be equipped with TEDs and hence a 2% average accuracy drop across various DNN datasets was reported.
Similar to EFFORT, in (Fan et al., 2022) a lightweight error detection circuit is introduced to detect timing errors from reduced voltage operations. However, circuit level techniques substantially complicate design and test process and are not applicable to already fabricated chips. Notice, that our proposed solution does not compromise the accuracy of DNN. MATIC (Kim et al., 2018) combines algorithm-level and circuit-level techniques to enable efficient low-voltage operation of neural accelerators. No online error detection is utilized, instead, MATIC addresses the challenge of timing errors based on speculative fault model and error-aware training. Hence, once the fault model no longer is applicable, the accuracy can not be ensured.
This work can be extended for training as well. The solution was only showcased on a GPU, while a similar approach can be applied to other DNN accelerators, e.g., Google TPUs. Due to software based nature of this work, Shavette, is more cost efficient compared to hardware based methods, e.g., Razor (Ernst et al., 2003).
6. Conclusion
This study demonstrates the existence of a significant voltage margin in the default settings of a commercial GPU and possibility of reliably removing it. In computations of DNN models, the margin can be safely optimized by incorporating ABFT schemes. The time and energy overheads from ABFT are low, e.g., around 3.5% and 1.1% respectively. Nearly a quarter of energy could be saved without sacrificing accuracy of the neural network. The results indicate a promising avenue for enhancing the energy efficiency of AI accelerators, without expensive HW techniques or having to compromise reliability.